OpenStudy (anonymous):

WILL FAN AND MEDAL PLESE HELP What is the solution to the equation below? Round your answer to the nearest decimal places. 2^(3x)=73 A x=2.06 B x=0.48 C x=0.05 D x=6.19

OpenStudy (anonymous):

\[2^{3x}=73\\ 3x=\frac{\ln(73)}{\ln(2)}\]